OpenStudy (aripotta):

i'm really feeling like an idiot. i completely forgot how to do this. can someone refresh my memory?

Parth (parthkohli):

\[\large b^{n/a}=\sqrt[a]{b^n}\]

Parth (parthkohli):

Yeah, but because that's a negative, I think you should put a \(\large \sqrt[3]{216^{-2}}\)

Parth (parthkohli):

Remember this guy?\[\large a^{-n} = \dfrac{1}{a^n}\]

Parth (parthkohli):

Yes, and now find the cube root of that thing.

OpenStudy (chihiroasleaf):

use this \[a ^{-n}=\frac{ 1 }{a ^{n} }\]

OpenStudy (agent0smith):

\[216^{\frac{ -2 }{ 3 }} = \frac{ 1 }{216^{\frac{ 2 }{ 3 }} } = \frac{ 1 }{ \sqrt[3]{216^{2}}} \]

OpenStudy (agent0smith):

Easier to think of it that way

Parth (parthkohli):

here:\[\sqrt[3]{\dfrac{a}{b}} = \dfrac{\sqrt[3]{a}}{\sqrt[3]{b}}\]I know I am smashing you with too many properties :-P

Parth (parthkohli):

An easier way to think about it is to find the cube root of the numerator and the denominator separately

OpenStudy (agent0smith):

\[\frac{ 1 }{216^{\frac{ 2 }{ 3 }} } = \frac{ 1 }{ \left( 216^\frac{ 1 }{ 3 } \right)^2} = \frac{ 1 }{ \sqrt[3]{216^{2}}}\]

Parth (parthkohli):

To make some stuff simpler,\[\dfrac{1}{\sqrt[3]{(6^3)^2}} = \dfrac{1}{\sqrt[3]{(6^2)^3}}\]

Parth (parthkohli):

Yeah 8-)

OpenStudy (aripotta):

thanks :)

OpenStudy (agent0smith):

Probably easier to find the cube root of 216, THEN square it, than square 216, then find it's cube root... both methods yield the same answer.